1. What is India's opportunity cost of producing one cell phone?

2.  What is the USA’s opportunity cost of producing one cell phone?

3. Which country has a "Comparative Advantage" at producing cell phones?

4. & 5. What will be the Maximum & Minimum ‘Price’ for a cell phone (in terms of tee shirts traded)? Minimum Price: ____ Maximum Price: ____

6. List one possible “Price” for cell phone, in terms of tee shirts traded, that would make BOTH India & USA better off: _____

7. (6 pts): 1) Below, draw the two “Production Possibilities Curve” - one for each country, India and USA! Label the axes of the Graphs. Show amounts (maximum Quantities produced of cell phones & tee shirts) along the axes of both graphs. 2)  Point to (or, circle) the good each country will sell (specialize at) on that country's PPC graph - on the axis where the max. quantity of that good is shown.   3) Show they both can move to a position or point OUTSIDE THEIR OWN curve at the “Price” you gave in question 6.  [Draw your 2 graphs neatly and legibly. #7 is Worth 6 POINTS!]

8. (1) Victoria works at her Hair & Nail salon. She can complete 20 nails appointments a day, or 4 hair color appointments a day. What is her "opportunity cost" of 1 hair color appointment? [Give a number answer.]

9. (2) Write a short example of an “opportunity cost” you sacrificed (gave up) when you made a personal choice, recently. Just a few sentences thanks!

India USA

Tee Shirts: 18 . 24

Cell phones . 9 . 3
Answer #1

Answering only first four parts are mandatory by HomeworkLib policy

India USA
T shirts (T) 18 24
Cell phone (C) 9 3

1) for india

To produce 9 C, give up 18 T

So to produce 1 C, give up 18/9 = 2 T

So opportunity cost of 1 C = 2 T shirts

2) for USA

To produce 3 C, give up 24 T

So opportunity cost of 1 C, = 24/3 = 8 T shirts

3) country with lower opportunity cost of production of C in terms of T, will have comparative advantage in cell phone

so India has comparative advantage in C,

( Bcoz only 2 units T are given up for every C in india , as against 8 T in USA)

4) the minimum price = 2 T

maximum price = 8 T